BRICS Notes Series, 2005

December 23, 2005

This document is also available as PostScript and DVI.

References

[NS-05-6] Abstract, PDF.

Anders Møller, editor. PLAN-X 2006 Informal Proceedings, (Charleston, South Carolina, January 14, 2006), December 2005. ii+92.

[NS-05-5] Abstract, PDF.

Patrick Cousot, Lisbeth Fajstrup, Eric Goubault, Maurice Herlihy, Kim G. Larsen, and Martin Raußen, editors. Preliminary Proceedings of the Workshop on Geometry and Topology in Concurrency, GETCO ’05, (San Francisco, California, USA, August 21, 2005), August 2005. vi+44.

[NS-05-4] Abstract, PDF.

Scott A. Smolka and Jiří Srba, editors. Preliminary Proceedings of the 7th International Workshop on Verification of Infinite-State Systems, INFINITY ’05, (San Francisco, USA, August 27, 2005), June 2005. vi+64 pp.

[NS-05-3] Abstract, PDF.

Luca Aceto and Andrew D. Gordon, editors. Short Contributions from the Workshop on Algebraic Process Calculi: The First Twenty Five Years and Beyond, PA ’05, (Bertinoro, Forlì, Italy, August 1–5, 2005), June 2005. vi+239 pp.

[NS-05-2] Abstract, PostScript, PDF, DVI.

Luca Aceto and Willem Jan Fokkink. The Quest for Equational Axiomatizations of Parallel Composition: Status and Open Problems. May 2005. 7 pp. To appear in a volume of the BRICS Notes Series devoted to the workshop “Algebraic Process Calculi: The First Twenty Five Years and Beyond”, August 1–5, 2005, University of Bologna Residential Center Bertinoro (Forlì), Italy.

[NS-05-1] Luca Aceto, Magnus Mar Halldorsson, and Anna Ingólfsdóttir. What is Theoretical Computer Science? April 2005. 13 pp.